Ross is six years older than twice his sister Holly's age. The difference in their ages is 18 years. How old is Holly?
Answer by lwsshak3(6490)

Ross is six years older than twice his sister Holly's age. The difference in their ages is 18 years. How old is Holly.
..
let x = Ross' age=2y +6
let y = Holly's age
x-y=18
y=x-18
x=2y+6
x=2(x-18)+6
x=2x-36+6
-x=-30
x=30
y=x-18=0-18=12
ans:
Hollys age = 12 years
Ross' age =2y+6=24+6=30 years